OpenStudy (anonymous):

evaluate log(4)256 , can someone show me how to do this

OpenStudy (anonymous):

\[\log_{4} 256\]

OpenStudy (freckles):

\[\log_4(256)=x \\ 4^x=256\]

OpenStudy (freckles):

4 to what power gives you 256?

OpenStudy (anonymous):

\[\log_{4} 256=x\] \[4^x=256=4^{?}\]

OpenStudy (anonymous):

4^4=256

OpenStudy (sweetburger):

you could do the change of base formula and then use a calculator unless your calculator is capable of having a different base

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

4 right?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

thats so simple thanks so much!!

OpenStudy (anonymous):

yw